OpenStudy (anonymous):

How do you divide polynomial when the exponent of the numerator is less than the denominator, for exmaple a^3/b^7?

OpenStudy (anonymous):

Sorry, I don't know what you mean.

OpenStudy (anonymous):

\[\frac{ a^3 }{ b^7 } \]

OpenStudy (anonymous):

Is the answer ab^4?

OpenStudy (loser66):

nope,

OpenStudy (anonymous):

a^4b?

OpenStudy (loser66):

you said that they are polynomial, so, only the one way you can do is factor them separately consider whether they have the same term or not, if they are cancel out those same term done. cannot make division.

OpenStudy (loser66):

give me a particular problem, cannot give out the clear explanation

OpenStudy (anonymous):

So, it is already simplified?

OpenStudy (loser66):

as you said above, a ^3 cannot divided by b^7

OpenStudy (loser66):

but if they are polynomial, it's another problem.

OpenStudy (loser66):

hey, it's not polynomial, friend

OpenStudy (loser66):

\[8=2^3\] therefore, hey, where is it?

OpenStudy (anonymous):

\[(\frac{ a^3 }{ b^7 })^2\]

OpenStudy (loser66):

tell me what you think?

OpenStudy (zzr0ck3r):

sorry my response was for a dif question...lol

OpenStudy (loser66):

ok, @zzr0ck3r you can continue now, appreciate

OpenStudy (anonymous):

First, multiply a^3 by 2 which is a^6 and then b^7 by 2 which is b^14 and then the answer is a^6/b^14?

OpenStudy (loser66):

yeap

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (loser66):

hey, why do you ask for polynomial, and then give out the example of an exponent of a constant?

OpenStudy (anonymous):

I don't know how to do that either.

OpenStudy (loser66):

they are different topics. somehow you can apply this stuff but in some cases only.

OpenStudy (anonymous):

Okay, i'll keep that in mind.

OpenStudy (loser66):

another thing is if a = m b, you can combine them. I don't want to confuse you, but there are so many rules in this field. That's why I ask you for a particular case. ok. good luck